Fermi National Laboratory


Accelerator Update for December 8, 2004 - December 10, 2004

Wednesday December 8
The day shift began with the Antiproton Source (Pbar) stacking, with Tevatron (TeV) and Recycler conducting studies, and with MiniBooNE taking beam.

Stacking was halted at 10:09 AM for about twenty minutes so that TeV studiers could analyze fake injections of antiprotons.

Operations put the TeV through a dry squeeze at 11:13 AM.

At 12:03 PM, FESS personnel checked out a sump problem in Enclosure C. They found a break in a discharge line. They installed a temporary pump.

At 12:18 PM, Pbar tripped off for the first of several times due to an Accumulator kicker problem. Experts investigated.

At 12:48 PM, a Linac RF station tripped off. Operators had the station back up by 1 PM.

Operations put the TeV into shot setup at 1:53 PM and established store 3824 at 2:50 PM. Its initial luminosity was 45.5E30.

At 5:05 PM, Operations reported that a TeV horizontal separator (D11) spark had caused a loss of luminosity.

Thursday December 9
The midnight shift began with Operations monitoring store 3824, with Pbar stacking, Recycler experts conducting studies, and with MiniBooNE taking beam.

At 12:19 AM, the TeV suffered a quench at sector E1 and lost the store. MIRF 11 tripped off just before the quench. Operations reset the station and turned it on at 12:41 AM, but it tripped off again and would not reset. They left the station off.

Operations reported at 2:16 AM that a TeV separator (D11) spark had caused the quench. Operators put all the separators into condition mode and then put the TeV through a dry squeeze.

Recycler experts ended their studies at 4:21 AM.

At 4:48 AM, FESS personnel began work on the temporary sump discharge line for Enclosure C. The existing discharge pipe must be replaced.

Operators began preparing the TeV and Main Injector (MI) for an access at 8:09 AM.

Operators had Linac and Booster prepared for an access by 8:49 AM.

A Booster expert reported at 9:57 AM that he found a LCW leak on a magnet hose for V802.

MI experts reported at 10:23 AM that magnet HV703 had a LCW leak.

The TeV came out of access and Operations began its checkout at 12:19 PM. Among other work, the SY magnet in TeV sector F4 had its bus reworked.

While conducting a QBS ramp check at 12:44 PM, the TeV tripped off due to a sector D4 lead failure. At 1:24 PM, experts decided they would have to access to investigate the D4 lead problem.

The Neutron Therapy Facility (NTF) began taking beam at 1:33 PM.

Operations reported that MiniBooNE, Pbar, and the Recycler were all taking beam by 3:39 PM. TeV turn on was just beginning.

Operations reported at 3:51 PM that techs completed the sump work in Enclosure C.

Operations put the TeV into shot setup at 6:14 PM, but a problem with the TeV abort put the setup on hold. After trying several things, operators got the abort to reset at 7:31 PM.

Operations established store 3829 at 9:22 PM with an initial luminosity of 47.93E30. Pbar resumed stacking. SY experts began trying to send beam to the SY dump.

Friday December 10
The midnight shift began with Operations monitoring store 3829, with Pbar stacking, with Recycler conducting studies, with SY experts working on their beam, and with MiniBooNE taking beam.

At 1:48 AM, the H- Source had a strange event where settings went to zero or to weird values. Everything reset okay.

At 6:16 AM, the MI power supply interlock bit dropped out for the MI-60 Quad PS. Operators called an expert who had the power supply repaired by 7:45 AM. He said he found a mouse in the trip module. The mouse, who had no good reason for being there, was being held for further questioning.

Pbar resumed stacking at 7:47 AM.

Plans for Friday and the Weekend
The plan for today includes a two-hour end-of-store study, beginning at 11 AM, followed by a four to eight hour Tevatron study. Shot setup will follow the studies. After the store is established, the Recycler will take shots of antiprotons from Pbar, which will probably occur early on the Saturday midnight shift.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/8/2004   email Fermilab